A room with a heat loss of 3 kW through walls and
Answer #1

heat loss = Q = 3kW = 3000 W

Heat transfer coefficient = U = 5 W K-1 m-2

Area = A = 6 m2

T1 = 20 0C

T2 = ?

Q = UA(T2-T1)

3000 = (5)(6)(T2-20)

T2-20 = 100

T2 = 100-20

= 120 oC

Correct temperature = (120 - 20) oC = 100 oC

Answer is D.